Astrid1:

@sNoWfLaKe0531 hElP mE aGaIn oR eLsHe.. *Uploading ss*

Astrid1:

snowflake0531:

\[m^2=14\] we need to get rid of that square \[\sqrt m^2 = \sqrt14\] simplifies to ... which one

Astrid1:

A?

snowflake0531:

i practically gave you the answer astrid ._.

Astrid1:

Oh- Sho itsh B?

snowflake0531:

yesh xd

Astrid1:

Ahhh. I'm stohhpihd. Now help me with the question below tht.

snowflake0531:

We have sqrt7 What squared is between 6.9 and 7.1 sqrt7 = 2.64575131106 First try 2.6 \[2.6^2 =?\]

Astrid1:

Uhhh 6.76?

snowflake0531:

Yep so is 6.76 between 6.9 and 7.1?

Astrid1:

Yesh. I tink. o:

snowflake0531:

can i die? are you sure that 6.76 is between 6.9 adn 7.1 ._____________.

Astrid1:

So it's not? Geehhzz, I'm soohptihd.

snowflake0531:

._. i suggest knowing the signs of < and > and knowing your decimals AnYwAyS So then we move on to the next decimal point 2.65 what is 2.65^2

Astrid1:

7.0225? o: I hope I am correct e.e

snowflake0531:

yes and is 7.0225 in between 6.9 and 7.1?

Astrid1:

Yes? (I'm probably wrong o-o)

snowflake0531:

._. yes, it is So AnYWaYs your answer is 2.65._.

Astrid1:

._. Thanks. I'll be back for more stupid asking 4 help. Jbc I'm stoopid.

snowflake0531:

lollllllllllllll it's fine xddddd and yw~